# MD-Splatting: Learning Metric Deformation from 4D Gaussians in Highly Deformable Scenes

### v1

@@ Abstract
Accurate 3D tracking in highly deformable scenes with occlusions and shadows can facilitate new applications in robotics, augmented reality, and generative AI. However, tracking under these conditions is extremely challenging due to the ambiguity that arises with large deformations, shadows, and occlusions. We introduce MD-Splatting, an approach for simultaneous 3D tracking and novel view synthesis, using video captures of a dynamic scene from various camera poses. MD-Splatting builds on recent advances in Gaussian splatting, a method that learns the properties of a large number of Gaussians for state-of-the-art and fast novel view synthesis. MD-Splatting learns a deformation function to project a set of Gaussians with non-metric, thus canonical, properties into metric space. The deformation function uses a neural-voxel encoding and a multilayer perceptron (MLP) to infer Gaussian position, rotation, and a shadow scalar. We enforce physics-inspired regularization terms based on local rigidity, conservation of momentum, and isometry, which leads to trajectories with smaller trajectory errors. MD-Splatting achieves high-quality 3D tracking on highly deformable scenes with shadows and occlusions. Compared to state-of-the-art, we improve 3D tracking by an average of 23.9 %, while simultaneously achieving high-quality novel view synthesis. With sufficient texture such as in scene 6, MD-Splatting achieves a median tracking error of 3.39 mm on a cloth of 1 $\times$ 1 meters in size. Project Website: https://md-splatting.github.io
Figure 1: We propose MD-Splatting, a method that improves state-of-the-art methods for simultaneous dynamic novel view synthesis and accurate 3D point tracking in highly deformable scenes. This figure shows the rendering and tracking of MD-Splatting in the six dynamic Blender [11] scenes used for evaluation. We will refer to the scenes in this Figure as Scenes 1, 2, 3, 4, 5 and 6 ordered from left to right.
@@ 1 Introduction
Three-dimensional tracking of clothing, towels, and similarly highly deformable objects as they drop, deform, and drape can facilitate new applications and improve the capabilities of existing applications. Generative AI and augmented reality applications can benefit from novel view synthesis for training, interaction, and scene editing [21]. Robots can benefit from tracking task-relevant geometries to dexterously manipulate deformable objects [46, 8, 47] and use novel-view synthesis to augment learning [20]. Simulation methods can improve fidelity using real2sim and related approaches that leverage real-world tracking data for parameter and method tuning [28]. However, tracking moving and morphing points poses many non-trivial challenges: points change color as they pass through shadows, occlusions interfere with the continuous visibility of points, and internal and external forces, such as material tension and air currents, cause points to move unpredictably.
Recent work demonstrated Gaussian splatting [25] can yield state-of-the-art novel view synthesis and rendering speeds exceeding 100 fps. 3D Gaussian Splatting fits the properties of a set of Gaussians, their color, position, and covariance matrix, using a fast differentiable rasterizer. An extension of 3D Gaussian splatting [33] showed dynamic scenes can be modeled by explicitly optimizing the properties of Gaussians over time, including the metric deformations. Metric deformations are the changes in the position of Gaussians in the world frame between two adjacent timesteps. Metric deformations across many timesteps yields 3D tracking of Gaussians, contrary to related work [42] with deformations available only in a canonical frame.
Explicitly optimizing the Gaussian pose as in Dynamic 3D Gaussians [33], the most relevant baseline to MD-Splatting, results in degraded performance with large deformations and shadows. Training time also scales linearly with the number of timesteps provided for each camera, meaning longer sequences take propotionally longer to train.
@@
We propose MD-Splatting, a method that uses time-synchronized image frames from a calibrated multi-camera setup to track 3D geometries of deformable objects as they move through shadows and occlusions. MD-Splatting learns the canonical state of a set of Gaussians and a deformation function that maps the Gaussians into metric space. This enables tracking by recovering metric deformations, and novel-view rendering (through splatting) using a fast differentiable rasterizer.
@@
We evaluate MD-Splatting in six photo-realistic synthetic scenes of varying difficulty. The scenes contain large deformations, shadows, and occlusions (see Figure 1 for visualization of the scenes and tracking trajectories by MD-Splatting). Empirical results show that MD-Splatting infers 16.7 % more accurate 3D tracking results compared to previous state-of-the-art [33], and achieves high-quality novel-view reconstruction results with an average PSNR of 39.1. In a scene with cameras at 1 m radius from the origin (i.e. Scene 6), MD-Splatting can track cloth deformation with as low as 3.39 mm median tracking error. To the best of our knowledge, MD-Splatting is the first approach for 3D tracking and view synthesis using canonical Gaussian splats.
In summary, we contribute:
  • The first approach that learns continuous metric deformations from a canonical space of Gaussians.
  • Experiments that suggest state-of-the-art performance in simultaneous 3D metric tracking and novel view synthesis. MD-Splatting improves tracking accuracy by an average of 16.7 % while achieving competitive view reconstruction.
@@
  • A set of six synthetic scenes with large deformations, strong shadows, and occlusions. The scenes will be shared with the community.
@@ 2 Related Work
MD-Splatting builds on prior work in novel view synthesis in dynamic scenes to enable simultaneous 3D tracking and novel view synthesis. We discuss prior works in novel view synthesis in static and dynamic scenes, as well as methods for point tracking.
@@ 2.1 Neural Rendering for Novel View Synthesis
MD-Splatting builds on prior work in novel view synthesis, and uses photometric consistency as a signal to achieve 3D tracking. A popular novel view synthesis approach is NeRF [34], which uses neural networks to learn scene representations that are capable of photo-realistic novel view reconstruction. The learned networks render a 3D point by outputting its RGB radiance and density when viewed from an arbitrary camera view. NeRF then renders pixels using existing volume rendering techniques. Subsequent works improve NeRF along several axes: one is to speed up training and inference time via novel representations and system optimizations [36, 44, 31, 68, 50, 19, 32, 36, 35], or depth supervision [12, 1, 63, 37, 49]. Other works extend NeRF to more challenging conditions, such as sparser camera views [69, 10, 54, 39], fewer extrinsic camera calibrations [29, 66, 9, 24], transparent objects [26, 23] and reflective surfaces [57].
Particle-based methods use a more explicit representation than typical NeRF-based approaches. MD-Splatting builds on 3D Gaussian Splatting [25], which proposed a differential rasterizer to render a large number of Gaussian ‘splats,’ each with their state including color, position, and covariance matrix. Gaussian splatting achieves real-time rendering of novel views with state-of-the-art performance.
@@ 2.2 Dynamic Novel View Synthesis
The assumption of static scenes in neural rendering approaches prevents application to real-world scenarios with moving objects and people, such as the dynamic and deformable scenes in this work. One line of work to address this assumption is adding a time dimension to NeRF modeling [15, 18, 27, 65], either condition the neural field on explicit time input or a time embedding. Another line of work learns a deformation field to map into a canonical space [43, 41], i.e., every 4D point in space and time maps to a 3D point in a canonical NeRF. This approach can be improved via distinguishing between foreground and background [52], or via leveraging depth priors [1, 63, 37], and can achieve fast training [16, 53].
To accelerate the training of dynamic neural radiance fields, DeVRF [30] proposed to model the 3D canonical space and 4D deformation field of a dynamic, non-rigid scene with explicit and discrete voxel-based representations. To solve the overfitting issue resulting from a large number of grid parameters, DeVRF [30] introduces a static-to-dynamic learning paradigm to first reconstruct a 3D canonical voxel space using static multi-view images and then optimize a 4D deformation voxel space using few-view dynamic sequences. It achieves convergence of optimizing dynamic voxel radiance fields within $10$ minutes on a single NVIDIA GeForce RTX3090 GPU. However, since it relies on the static scene training stage to learn the scene information, it struggles to reconstruct dynamic scenes with complex deformations and shadows, especially with dynamic occlusions.
Several recent works extend the above approaches to 3D Gaussian splatting. Dynamic 3D Gaussians [33] explicitly model the position and covariance matrix of each Gaussian at each time step. This method struggles in dynamic scenes with large deformations, strong shadows, or occlusions. We build on another recent work, 4D Gaussian splatting [64], which uses feature encoding techniques proposed in HexPlanes [6] and K-planes [17], and learns a deformation field instead. To the best of our knowledge, all prior deformation-based work learned deformations in a non-metric canonical space without reconstructing metric deformations. We propose a method to learn metric deformations from a set of Gaussians with canonical properties, deformed to metric space using a neural-voxel deformation function. By training on adjacent timesteps, MD-Splatting allows enforcing physics-informed regularization terms to improve performance beyond photometric consistency.
@@ 2.3 Point Tracking
Point tracking methods, usually trained on large amounts of data, aided previous 3D tracking approaches by providing a strong prior [30]. We also construct several baselines that include point tracking methods (Section 6). Prior work on point tracking often studies tracking 2D points across video frames, where a dominant approach is training models on large-scale synthetic datasets containing ground-truth point trajectories [13, 14, 71, 22] or dense optical flows [60]. Optical flow [3, 51] or scene flow [56, 2, 58] can also be viewed as single-step point-tracking in 2D and 3D, respectively.
@@
A main motivation for studying point tracking is the potential it can unlock for robotics applications: for example, RoboTAP [55] shows pre-trained point-tracking models improves sample efficiency of visual imitation learning.
@@ 3 Problem Statement
3D Tracking The 3D tracking objective is to recover the trajectory of any point in a dynamic scene by modeling metric deformations. Thus, the objective is to find a function $x_{t}=Q(x_{0},t_{0},t)$, where $x_{0}\in\mathbb{R}^{3}$ is the location of a point of interest at a chosen time $t_{0}$, while $x_{t}\in\mathbb{R}^{3}$ is the location of the same point at another chosen time $t$. The function $Q$ is valid for any point $x_{0}$ and any $t\in[0,H]$, allowing for tracking of any point in space.
Novel View Synthesis This objective is to recover novel views from arbitrary viewpoints. The extrinsics at any viewpoint can be captured by matrix $P$, with $P=K[R|T]$. Here $K$ is the intrinsics matrix, $R$ is the rotation matrix of a camera with respect to the world frame, and $T$ is the translation vector with respect to the world frame. The aim is to learn a function $V$ such that $I_{P,t}=V(P,t)$, where $I_{P,t}$ is an image rendered from a camera with extrinsics $P$ at time $t$. As with the tracking objective, the time parameter is valid for any $t\in[0,H]$.
@@ 4 Preliminary
4.1 Neural Radiance Fields (NeRF)
Neural radiance fields (NeRFs) trace rays through multiple camera views to train an MLP $\Phi_{r}$ to infer a radiance (RGB) and density ($\sigma)$ from a world coordinate $(x)$ and view direction $(\theta,\phi)$ inputs, which are typically concatenated with a positional encoding in the form of sinusoids with varying frequencies.
$\Phi_{r}(x,\theta,\phi)=(RGB\sigma)$ (1)
@@ 4.2 Gaussian Splatting
3D Gaussian Splatting [25] deploys a more explicit scene representation than NeRFs, by rendering a large set of Gaussians each defined by their mean position $\mu$ and covariance matrix $\Sigma$. Thus, for each $x\in\mathbb{R}^{3}$, its Gaussian $G(x)$ is
@@
Directly optimizing the covariance matrix $\Sigma$ would lead to infeasible covariance matrices, as they have to be positive semi-definite to have a physical meaning. Instead, Gaussian Splatting [25] proposes to decompose $\Sigma$ into a rotation $R$ and scale $S$ for each Gaussian:
@@
where $c_{i}$ is the color of each point and $\alpha_{i}$ is given by evaluating a 2D Gaussian with covariance multiplied with a learned per-point opacity [67, 25].
@@ 4.3 Deformation Fields for Dynamic Scenes
Prior work showed that a deformation function combined with a static NeRF in a canonical space can enable novel view synthesis in dynamic scenes. The deformation function $F_{\mathrm{NeRF}}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ deforms a point in world coordinates $(x^{\prime})$ into a point in canonical space $(x)$. Prior work formulated $F_{\mathrm{NeRF}}$ as an MLP [42] and a multi-resolution voxel grid [16]. Wu et al. [64] applied a similar approach to arrive at Gaussian splatting of dynamic scenes. Given the state of a single canonical Gaussian, defined by $P=[\mu,S,R,\sigma,C]$ at time $t$, a deformation function takes the form
@@
We propose MD-Splatting (Figure 2), based on 4D-Gaussians [64], to render novel views in dynamic scenes. The key differences with 4D-GS are: (1) we propose a simple method to track canonical Gaussians in 3D using a canonical $\rightarrow$ metric deformation function, (2) the output of the deformation function is different, and (3) using the method shown in Figure 3, we enforce regularization losses on the 3D trajectories of Gaussians.
Figure 2: MD-Splatting maps a set of Gausians with canonical properties to metric space using a deformation function $F$. The deformation function takes in the position of a Gaussian $(x,y,z)$ and a queried timestamp $t$, to infer shadow $s$, rotation $r^{\prime}$ and metric position $x^{\prime}$. During training, we use the metric positions and rotations to regularize the deformation function, considering the state at $t=\{i-1,i,i+1\}$ with Gaussian metric states $P^{\prime}_{t-1},P^{\prime}_{t},P^{\prime}_{t+1}$

### v2

# DeformGS: Scene Flow in Highly Deformable Scenes for Deformable Object Manipulation

@@ Abstract
Teaching robots to fold, drape, or reposition deformable objects such as cloth will unlock a variety of automation applications. While remarkable progress has been made for rigid object manipulation, manipulating deformable objects poses unique challenges, including frequent occlusions, infinite-dimensional state spaces and complex dynamics. Just as object pose estimation and tracking have aided robots for rigid manipulation, dense 3D tracking (scene flow) of highly deformable objects will enable new applications in robotics while aiding existing approaches, such as imitation learning or creating digital twins with real2sim transfer. We propose DeformGS, an approach to recover scene flow in highly deformable scenes, using simultaneous video captures of a dynamic scene from multiple cameras. DeformGS builds on recent advances in Gaussian splatting, a method that learns the properties of a large number of Gaussians for state-of-the-art and fast novel-view synthesis. DeformGS learns a deformation function to project a set of Gaussians with canonical properties into world space. The deformation function uses a neural-voxel encoding and a multilayer perceptron (MLP) to infer Gaussian position, rotation, and a shadow scalar. We enforce physics-inspired regularization terms based on conservation of momentum and isometry, which leads to trajectories with smaller trajectory errors. We also leverage existing foundation models SAM and XMEM to produce noisy masks, and learn a per-Gaussian mask for better physics-inspired regularization. DeformGS achieves high-quality 3D tracking on highly deformable scenes with shadows and occlusions. In experiments, DeformGS improves 3D tracking by an average of 55.8 % compared to the state-of-the-art. With sufficient texture, DeformGS achieves a median tracking error of 3.3 mm on a cloth of 1.5 $\times$ 1.5 m in area. Website: https://deformgs.github.io
Keywords:
Perception, Machine Learning in Robotics , Manipulation & Grasping
@@ 1 Introduction
Figure 1: We propose DeformGS, a method that improves state-of-the-art methods for accurate 3D point tracking in highly deformable scenes. This figure shows the rendering and tracking of DeformGS in the six dynamic Blender [9] scenes used for evaluation. We will refer to the scenes in this Figure as Scenes 1, 2, 3, 4, 5 and 6 ordered from left to right.
Recent advances in robot learning have demonstrated impressive performance on challenging tasks, including rigid and deformable object manipulation. Scaling these approaches to deployment will require an improvement in robustness and learning from few demonstrations. A promising avenue for improving in robot learning performance are intermediate representations and foundation models, including 6D object pose estimation [10, 11, 27, 42, 50, 51, 55], semantic latent features [35], and 2D pixel-wise tracking [21, 52]. However, perception and representations that will lead to robust manipulation of deformable objects remains an open challenge, due to self-occlusions, shadows, and varying (or lack of) textures.
Three-dimensional dense point tracking, or 3D scene flow, is a useful representation for robot manipulation, as it provides flexibility to represent high-dimensional dynamic state changes, while the deformable objects drops, deforms, and drapes during manipulation. In particular, dense 3D scene flow can be an input to imitation learning policies [52, 3], can be used to learn a transition model [40], to identify and track task-relevant key points, or to create a digital twin through real2sim transfer. Recent work in monocular tracking has seen improvements in performance on datasets such as TAP-Vid [13], but it remains unclear how to effectively lift from 2D tracking to 3D for robotic spatial understanding in challenging highly deformable scenes.
To overcome these limitations, Gaussian Splatting provides a promising avenue. Recent work demonstrated Gaussian Splatting [22, 23] can yield state-of-the-art novel-view synthesis and rendering speeds exceeding 100 fps. Concretely, 3D Gaussian Splatting uses a fast differentiable renderer to fit the colors, positions, and covariances of a set of Gaussians. An extension of 3D Gaussian splatting [29] showed dynamic scenes can be modeled by explicitly optimizing the properties of Gaussians over time, resulting in novel-view synthesis and scene flow.
Explicitly optimizing the Gaussian pose as in Dynamic 3D Gaussians [29] may result in degraded performance with large deformations and shadows. The Gaussian properties may converge to local optima, especially in scenes with large deformations, strong shadows, and occlusions.
@@
We propose DeformGS, a method that uses time-synchronized image frames from a calibrated multi-camera setup to track 3D geometries of deformable objects as they move through shadows and occlusions. DeformGS learns the canonical state of a set of Gaussians and a deformation function that maps the Gaussians into world space. This enables tracking by recovering scene flow, and novel-view rendering (through splatting) using a fast differentiable rasterizer.
@@
We evaluate DeformGS in six photo-realistic synthetic scenes of varying difficulty. The scenes contain large deformations, shadows, and occlusions (Figure 1 shows the scenes and tracking trajectories computed by DeformGS). Empirical results show that DeformGS infers 55.8 % more accurate 3D tracking results compared to previous state-of-the-art [53, 29]. In a scene with a 1.5 m $\times$ 1.5 m cloth (i.e., Scene 1), DeformGS can track cloth deformation with as low as 3.3 mm median tracking error.
We also evaluate DeformGS in the real world on the Robo360 [26] dataset. We show qualitative results for tracking rigid and deformable objects in cluttered scenes, and study two robotics applications: (1) real2sim transfer to create a digital twin, and (2) tracking task-relevant keypoints for downstream grasping applications.
In summary, our contributions are as follows:
  • We provide the first approach designed to accurately perform 3D dense tracking for deformable objects using 4D Gaussians.
  • We provide experiments that suggest state-of-the-art performance in simultaneous 3D metric tracking and novel view synthesis. DeformGS improves tracking accuracy by an average of 55.8% in synthetic experiments and demonstrates robust 3D tracking in the real world for deformable objects. The latter can be exploited as a representation for imitation learning and represents a new method for building digital twins.
@@
  • A set of six synthetic scenes with large deformations, strong shadows, and occlusions. We will open-source the scenes and as well as the source code.
@@ 2.1 Neural Rendering for Novel View Synthesis
DeformGS builds on prior work in novel-view synthesis, and uses photometric consistency as a signal to achieve 3D tracking. A popular novel view synthesis approach is NeRF [30]. It uses neural networks to learn scene representations that are capable of photo-realistic novel view reconstruction. Particle-based methods use a more explicit representation than typical NeRF-based approaches. DeformGS builds on 3D Gaussian Splatting [22, 23] which belongs to the latter caregory. [22] proposed a differential rasterizer to render a large number of Gaussian ‘splats,’ each with their state including color, position, and covariance matrix. Contrary to the NeRF-based approaches, Gaussian splatting achieves real-time rendering of novel views with state-of-the-art performance.
@@ 2.2 Dynamic Novel View Synthesis
The assumption of static scenes in neural rendering approaches prevents application to real-world scenarios with moving objects or humans, such as the dynamic and deformable scenes in this work. One line of work to address this assumption is adding a time dimension to NeRF modeling [15, 18, 25, 54]. Prior works either condition the neural field on explicit time input or a time embedding. Another line of work learns a deformation field to map 4D points into a canonical space [37, 36], i.e., every 4D point in space and time maps to a 3D point in a canonical NeRF. DeVRF [28] proposed to model the 3D canonical space and 4D deformation field of a dynamic, non-rigid scene with explicit and discrete voxel-based representations.
Several recent works extend the above approaches to 3D Gaussian splatting. Dynamic 3D Gaussians [29] explicitly model the position and covariance matrix of each Gaussian at each time step. This method struggles in dynamic scenes with large deformations, strong shadows, or occlusions. We build on another recent work, 4D Gaussian splatting [53], which uses feature encoding techniques proposed in HexPlanes [6] and K-planes [17], and learns a deformation field instead.
@@ 2.3 Point Tracking
Point tracking methods, usually trained on large amounts of data, aided previous 3D tracking approaches by providing a strong prior [28]. We also construct several baselines that include point tracking methods (Section 6). Prior work on point tracking often studies tracking 2D points across video frames, where a dominant approach is training models on large-scale synthetic datasets containing ground-truth point trajectories [12, 14, 57, 20] or dense optical flows [48]. Optical flow [2, 43] or scene flow [45, 1, 46, 19] can also be viewed as single-step point-tracking in 2D and 3D, respectively.
@@
2.4 Tracking for Robotics
A core motivation for studying point tracking is the potential it can unlock for robotics applications: for example, RoboTAP [44] shows pre-trained point-tracking models improve sample efficiency of visual imitation learning. It detects task-relevant keypoints, infers where those points should move to, and computes an action that moves them there. Any-point [52] learns to predict keypoint tracks, but conditioned on language inputs. Track2Act [3] builds on Any-point by learning a generalizable zero-shot policy, which only needs a few embodiment-specific demonstrations.
Rigid-body, or 6D, pose tracking and estimation has a rich history in robotics due to it foundational ability to model the world for a robot to manipulate [49, 31, 10, 11, 27, 42, 50, 51, 55]. In this work, we propose a deformable object analog of 6D pose tracking with the aim of extending successes to deformable object manipulation.
While existing methods leverage 2D tracking, and learn an additional policy to output robot actions, DeformGS provides a more powerful representation that allows for reasoning directly in 3D, instead of in the 2D image space.
@@ 3 Problem Statement
3D Tracking The primary goal is to recover the trajectory of any point in a dynamic scene by modeling the deformation of Gaussians over time. Thus, the objective is to find a function $x_{t}=Q(x_{0},t_{0},t)$, where $x_{0}\in\mathbb{R}^{3}$ is the location of a point of interest at a chosen time $t_{0}\in[0,H]$, while $x_{t}\in\mathbb{R}^{3}$ is the location of the same point at another chosen time $t\in[0,H]$. The function $Q$ is valid for any point $x_{0}$ and any $t\in[0,H]$, allowing for tracking of any point in space.
Novel View Synthesis The secondary goal is to achieve accurate scene flow by using photometric consistency as a supervision signal. To achieve this, the objective is to recover novel views from arbitrary viewpoints. The extrinsics at any viewpoint can be captured by matrix $P$, with $P=K[R|T]$. Here $K$ is the intrinsics matrix, $R$ is the rotation matrix of a camera with respect to the world frame, and $T$ is the translation vector with respect to the world frame. Concretely, the goal is to learn a function $V$ such that $I_{P,t}=V(P,t)$, where $I_{P,t}$ is an image rendered from a camera with extrinsics $P$ at time $t$. As with the tracking objective, the time parameter is valid for any $t\in[0,H]$.
@@ 4.1 Gaussian Splatting
3D Gaussian Splatting [22] deploys an explicit scene representation by rendering a large set of Gaussians each defined by their mean position $\mu$ and covariance matrix $\Sigma$. Given $x\in\mathbb{R}^{3}$, its Gaussian is
@@
Directly optimizing the covariance matrix $\Sigma$ would lead to infeasible covariance matrices, as they must be positive semi-definite to have a physical meaning. Instead, Gaussian Splatting [22] proposes to decompose $\Sigma$ into a rotation $R$ and scale $S$ for each Gaussian:
@@
where $c_{i}$ is the color of each Gaussian and $\alpha_{i}$ is given by evaluating a 2D Gaussian with covariance multiplied with a learned per-Gaussian opacity $\sigma$ [56, 22]. This representation allows for fast rendering of novel views, and aims to reconstruct the geometry of the scene.
@@ 4.2 Deformation Fields for Dynamic Scenes
Prior work showed that a deformation function combined with a static NeRF in a canonical space can enable novel view synthesis in dynamic scenes. The deformation function $F_{\mathrm{NeRF}}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ deforms a point in world coordinates $(x^{\prime})$ into a point in canonical space $(x)$.
Prior work formulated $F_{\mathrm{NeRF}}$ as an MLP [38] and a multi-resolution voxel grid [16]. Wu et al. [53] applied a similar approach to arrive at Gaussian splatting of dynamic scenes. Given the state of a single canonical Gaussian, defined by $P=[\mu,S,R,\sigma,C]$ at time $t$, a deformation function is
@@
We propose DeformGS (Figure 2), based on 4D-Gaussians [53], to render novel views in dynamic scenes. The key differences with 4D-GS are: (1) we propose an intuitive method to track canonical Gaussians in world coordinates using a continuous deformation function, (2) the output of the deformation function is different, e.g., DeformGS infers shadows and does not alter opacity or scale over time, and (3) using the method shown in Figure 3, we enforce physics-inspired regularization losses on the 3D trajectories of Gaussians.
Figure 2: DeformGS maps a set of Gausians with canonical properties to metric space using a deformation function $F$. The deformation function takes in the position of a Gaussian $(x,y,z)$ and a queried timestamp $t$, to infer shadow $s$, rotation $r^{\prime}$ and metric position $x^{\prime}$. During training, we use the metric positions and rotations to regularize the deformation function, considering the state at $t=\{i-1,i,i+1\}$ with Gaussian metric states $P^{\prime}_{t-1},P^{\prime}_{t},P^{\prime}_{t+1}$